# Trustworthiness of Stochastic Gradient Descent in Distributed Learning

### v1

@@ 5 Discussion
However, we observe that in certain configurations, gradient compression does show more noticeable effects. For instance, in the CIFAR-100 dataset, the use of Top-K SGD with higher compression levels (e.g., Level 30 and Level 50) leads to a more pronounced decrease in the attack success rates. This indicates that specific compression strategies and higher levels of compression can have a tangible impact on MIA resistance, although this effect is not consistently observed across all datasets and algorithms.
@@
These findings suggest that the relationship between gradient compression and MIA success rates may be more complex than initially assumed. While black-box MIA techniques may be unreliable in detecting subtle privacy leakage, it is also possible that certain gradient compression methods retain sufficient information for adversaries to exploit. This aligns with recent studies that argue current privacy evaluation metrics: MIAs may be misleading [27, 28].
6 Conclusion
@@
This study examines the impact of gradient compression techniques, including PowerSGD and Top-K SGD, and uncompressed SGD on the effectiveness of GradInv and MIA in distributed learning across various datasets. Unlike previous work, we provide comprehensive empirical evaluations to assess the privacy risks of different SGD variants, addressing gaps in the understanding of these methods’ trustworthiness.
@@
Our findings show that uncompressed gradients using Original SGD pose the highest privacy risk, as demonstrated by high SSIM values and detailed reconstructions. In contrast, gradient compression techniques significantly lower the quality of reconstructed images, thereby enhancing privacy protection. Notably, we find that MIA demonstrates low sensitivity to both compressed and uncompressed SGD, indicating that MIA may not be a reliable metric for assessing privacy risks in machine learning. We call for further investigation into the existing privacy evaluation methods for machine learning.
@@
Future work will explore the effects of compression on other attack vectors, such as data poisoning and backdoor attacks, to gain a more complete understanding of the privacy implications of these methods. Additionally, we aim to build a stronger theoretical foundation for assessing the trustworthiness of distributed learning algorithms.

### v3

@@ V Discussion
However, in certain configurations—particularly for CIFAR-100—Top-K SGD with higher compression levels (e.g., Level 1 and Level 2) shows a more pronounced decrease in attack success. This indicates that specific compression strategies at certain intensities can have a tangible impact on MIA resistance, though this pattern is not universal across all datasets and algorithms.
@@
Overall, these observations highlight a more complex relationship between gradient compression and MIA success than initially anticipated. On the one hand, gradient compression may preserve enough information for adversaries to exploit. On the other hand, current black-box MIA techniques may be insufficiently sensitive to subtle privacy leakages in machine learning, especially in distributed learning. This aligns with recent findings suggesting that existing MIA metrics can be misleading in assessing privacy risks [28, 29].
VI Conclusion
@@
This study examines the impact of gradient compression techniques, including PowerSGD and Top-K SGD, and uncompressed SGD on the effectiveness of GradInv and MIA in distributed learning across various datasets. We provide comprehensive empirical evaluations to assess the privacy risks of different SGD variants, addressing gaps in the understanding of these methods’ trustworthiness.
@@
Our findings show that uncompressed gradients using Original SGD pose the highest privacy risk, as demonstrated by high SSIM values and detailed reconstructions. In contrast, gradient compression techniques significantly lower the quality of reconstructed images, thereby enhancing privacy protection. Notably, we find that MIA demonstrates low sensitivity to both compressed and uncompressed SGD, indicating that MIA may not be a reliable metric for assessing privacy risks in distributed machine learning. We call for further investigation into the existing privacy evaluation methods for machine learning.
@@
Future work will explore the effects of compression on other attack methods, such as data poisoning and backdoor attacks, to gain a more complete understanding of the privacy implications of these methods.